Which of the following statements is true about spectrophotometry?

a. absorbance equals transmittance

b. solution with absorbance value of 0.8 absorbs twice as much as one with absorbance value of 0.4

c. transmittance is measured on a log scale

d. absorbance will be measured at 595nm in this lab

e. absorbance and transmittance are inversely related
Answer #1

Absorbance is inversely proportional to transmittance in spectrophotometry. When light is absorbed, it will not be transmitted by the solution. If light is not absorbed by a solution, it will be transmitted.

Absorbance of solution (A)= 2- log (% Transmittance of solution)

Suppose % transmittance= 100

Then A= 2- log (100)= 2-2= 0

Transmittance is given as percentage and not given on log scale.

As log of transmittance is considered for calculating absorbance, absorbance of 0.8 will not absorb as much as absorbance of 0.4. Whether the experiment is using 595 nm cannot be determined as the details of experiment are not included. However, absorbance measurement depends on the color of the solution and what filter is being used.

Right choice: e. absorbance and transmittance are inversely related.